OpenStudy (anonymous):

Shea and her best friend Kelly both babysit during the summer. Shea charges $4 per hour plus a $3 service fee, while Kelly charges $3.50 per hour plus a $6 service fee. Solve the system by substitution to find the number of hours, h, and the amount of pay, p, where they each earn the same amount.

OpenStudy (anonymous):

Help Please !

OpenStudy (anonymous):

4h+3=The Pay of Shea 3.5h +6= The pay of Kelly 4h+3=3.5h+6 -3 -3 4h=3.5h+3 -3.5h .5h=3 *2 *2 h=? Then stick "h" into one of the original equations and solve for someones pay to get "p" 3.5(h)+6=p OR 4(h)+3 =p